# Overcoming data challenges through enriched validation and targeted sampling to measure whole-person health in electronic health records

### v4

@@ 3.2.2 Wave II Chart Reviews
In Wave II, $3867$ data points were validated ($718$ labs, $3149$ vitals), which was fewer than in the Pilot + Wave I. This difference was explained by fewer patient encounters, with $16$-$453$ validated data points per patient (median $=65.5$). The estimated error rates for non-missing components (TPR $=99\%$, FPR $=3\%$) and missing data recovery rate ($17\%$) in Wave II closely resembled the Pilot + Wave I (Figure 6B). Still, some auditors’ findings differed (Figure 5B). For example, more vitals matched Epic ($65\%$).
Combined findings from all validation waves indicated that non-missing EHR data were of excellent quality (TPR $=99.5\%$, FPR $=2\%$), and auxiliary information could be located for $21\%$ of missing components (Figures 5C and 6C). On average, the number of non-missing components per patient increased from $6$ to $7$. Supplemental Tables S1–S2 contain detailed findings.
@@ 3.3 Analysis Findings: Whole-Person Health and Healthcare Utilization
Using ALI values from the EHR data, we fit first the naive healthcare utilization model. The expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 00 were $0.25$ (95% confidence interval [95% CI]: $0.18$, $0.36$). For every $0.1$-point (one-component) increase in ALI, these odds were expected to increase by $10\%$ (OR $=1.10$, 95% CI: $1.03$, $1.18$), adjusting for age. For every $10$ years older a patient was, their odds were expected to increase by $11\%$ (OR $=1.11$, 95% CI: $1.00,1.23$), adjusting for ALI. While these findings align with the literature, we wanted to ensure that missingness and potential errors in the EHR data did not lead to spurious conclusions.
3.3.1 Preliminary Estimates Based on the Pilot + Wave I Validation
Incorporating the Pilot + Wave I, we re-fit the healthcare utilization model with the first $52$ patients’ validated data and all $1000$ patients’ unvalidated data. These “preliminary SMLEs” were used to design the simulations in Section 2.6.1. Here, we compare them to the naive model. For an 18-year-old patient with an ALI of 00, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$). The association between whole-person health and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase in ALI (OR $=1.21$, 95% CI: $1.17,1.25$), adjusting for age. The adjusted OR $=1.10$ for age was relatively unchanged (95% CI: $1.01,1.21$). Thus, the preliminary SMLEs reaffirmed that worse whole-person health was associated with higher expected odds of engaging in the healthcare system, perhaps more strongly than the naive model indicated.
@@ 3.3.2 Final Estimates Based on All Waves of Validation
After completing all three waves, we used the $100$ patients’ validated data with all $1000$ patients’ unvalidated data to re-fit the model one last time. All final estimates fell between the naive model and preliminary SMLEs. The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 00 were $0.23$ (95% CI: $0.16$, $0.33$). For each $0.1$-point increase in ALI, a $12\%$ increase in odds was expected (OR $=1.12$, 95% CI: $1.09,1.15$), adjusting for age. Again, older patients were expected to be more likely to engage, controlling for ALI (OR $=1.11$, 95% CI: $1.02,1.20$). Additional technical details, including R code, are available in the Supplemental Materials.
@@ 3.4 Informing Implementation in the Learning Health System
Ultimately, all models agreed that worse whole-person health (i.e., higher ALI) was associated with a higher likelihood of engaging in the healthcare system, adjusting for age. Chart reviews uncovered very high quality of non-missing EHR data and located auxiliary information for some missing data. Without validation, knowing whether we could trust the EHR data and whether the EHR-derived ALI was still a valid computable phenotype for whole-person health would not have been possible. Thus, while the models led to the same conclusions here, validating EHR data before using them in practice or research is critical. Our findings can now inform implementation strategies to incorporate the ALI into this learning health system.
Having established ALI as a strong predictor of healthcare utilization, embedding it in the EHR is a critical next step. If ALI automatically prepopulated in patients’ electronic charts, clinicians would have easy access to it when making care decisions and researchers could extract it for further studies. This healthcare system previously implemented another score, the EFI, across the patient population aged $65+$. Building on that experience and encouraged by the ALI’s predictive ability, we plan to operationalize it for patients $18$ years and older.
Of course, we want to embed the validated ALI, and only a small subset of patients can undergo chart reviews. Fortunately, in obtaining SMLEs for the healthcare utilization model, we robustly estimated the exposure error mechanism, $\Pr(X|X^{*},Z)$. When scaling up across the healthcare system, the estimated exposure error mechanism can be used to efficiently predict patients’ validated ALI from unvalidated EHR data (Figure 7). See the Supplemental Materials for details on calculating these predictions.
Figure 7: Among patients with chart reviews (triangles), validated allostatic load index (ALI) could differ noticeably from the unvalidated version using the EHR data. For patients without chart reviews (squares), validated ALI could be robustly predicted using the estimated exposure error mechanism. The dashed line indicates the line of equality (i.e., where unvalidated and validated ALI are the same). The solid line is a loess smoother.
@@ 4 Discussion
Predicting healthcare utilization is a priority in the learning health system, with important implications at both patient and institutional levels. Establishing standardized, valid measures of whole-person health (like ALI) that are associated with utilization and can scale up to the entire institution is an integral step. While EHR data contain variables to calculate these measures, missing data is a key challenge, and non-missing data are still prone to errors.
@@
Data validation, often through chart reviews, can gauge data quality, but this time-intensive process is usually only possible for a small subset of patients. Also, existing validation protocols do not overcome missingness. We propose novel methods to maximize information gained from a partial validation study of $100$ patients from an EHR dataset of $1000$ through the statistical model, validation protocol, and targeted strategy to select patients. We focused on estimating the association between ALI and healthcare utilization. Partial validation can also efficiently improve estimates of the ALI itself, in preparation for implementation in the EHR.
@@
The proposed analytical and design strategies were devised to ensure robust and efficient estimates. Specifically, the statistical model avoids making restrictive assumptions about the error mechanism, relating the unvalidated and validated ALI, while incorporating all available information to achieve precise estimates (i.e., low variance). Moreover, the targeted design strategies are tailored to the preliminary validation data and chosen statistical model, selecting patients most informative for the clinical question. Our enriched validation protocol builds on the literature (e.g., [52]) by incorporating a clinically-driven “roadmap” to locate auxiliary information about missing values.
@@
The robust SMLEs can handle conditioning on error-prone $X^{*}$ and maybe one other covariate $Z$ in the error mechanism, due to the curse of dimensionality. We assumed that validated ALI depended only on its unvalidated version. However, if a more complicated error structure is expected, data pre-processing will be needed before using the SMLEs Tao2021.
@@
While our roadmap helped recover some missing ALI components, validated patients could still have incomplete data. Clinical guidelines were only defined for missing values presumed to be unhealthy. If auditors were given additional guidance on what auxiliary information might mean that a missing value is healthy, missingness could be further reduced. Others have described missingness in the EHR as selection bias and sought to correct for it using approaches like IPW [53]. Applying such approaches to our post-validation data may reduce the impact of residual missingness. Alternatively, we could adopt other allostatic load calculations, like the index of cardiometabolic health (ICMH), which rely less on rare labs and may face fewer missing components [9].
@@
EHR data are susceptible to “informative presence bias,” wherein the process through which patients engage in the healthcare system can bias associations [54, 55, 56]. Our decision to adapt the Seeman et al. [7] definition of the ALI as the proportion of non-missing components could be considered a summary measure. A recent review found summary measures to be the most common approach to handling informative presence bias [57].
@@
Whole-person health scores like ALI are among many computable phenotypes currently applied to EHR data [16]. Other phenotyping algorithms, including EHRshot [58], also seek to translate observable characteristics into measures of underlying patient states. Validation is a critical step in their development and typically involves chart reviews to obtain some “gold standard” labels [59, 60]. When phenotyping with machine learning methods, this validation set generally needs to be an SRS. Meanwhile, our statistical modeling approach accommodates targeted sampling of the validation study. This flexibility creates opportunities to promote the efficiency of downstream analyses (e.g., associations between the phenotype and suspected risk factors).
@@
There are several interesting directions for future work. First, a longitudinal approach to modeling healthcare utilization, wherein patients’ ALI can vary over time, might provide more information. Second, auditors recorded their findings in a Microsoft Excel workbook. For ease of use and accuracy, an electronic REDCap [61] interface could be built following published guidelines [35]. Third, targeted validation strategies could be used to improve predictive ability, rather than efficiency [62, 63].
@@ 5 Conclusion
EHR data provide an exciting opportunity to implement key measures in the learning health system. However, errors and missingness in these data pose challenges in verifying and operationalizing such measures. Validation of even a subset of patients’ data is time- and cost-intensive, but provides crucial insights into data quality. Statistical and biomedical informatics methods can maximize information gained from such partial validation.

### v5

@@ 3.2.2 Wave II Chart Reviews
In Wave II, $3867$ data points were validated ($718$ labs, $3149$ vitals), which was fewer than in the Pilot and Wave I. This difference was explained by fewer patient encounters, with 1616453453 validated data points per patient (median =65.5=65.5). The estimated error rates for non-missing components (TPR $99\%$, FPR $3\%$) and missing data recovery rate ($17\%$) in Wave II closely resembled the Pilot and Wave I (Figure 6B). Still, some auditors’ findings differed (Figure 5B), like more vitals matching Epic ($65\%$).
Combined findings from all validation waves indicated that non-missing EHR data were of excellent quality (TPR $99.5\%$, FPR $2\%$), and auxiliary information was located for $21\%$ of missing components (Figures 5C and 6C). The median number of non-missing components per patient increased from $6$ to $7$. Supplemental Tables S1–S2 contain detailed findings.
@@ 3.3 Analysis Findings: Whole-Person Health and Healthcare Utilization
Using the unvalidated EHR data, we first fit the naive model. The expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 0 were $0.25$ (95% confidence interval [95% CI]: $0.18$, $0.36$). For every $0.1$-point (one-component) increase in ALI, these odds were expected to increase by $10\%$ (OR =1.10=1.10, 95% CI: $1.03$, $1.18$), adjusting for age. For every $10$ years older a patient was, their odds were expected to increase by $11\%$ (OR =1.11=1.11, 95% CI: $1.00,1.23$), adjusting for ALI. While these findings align with the literature, we wanted to ensure that missingness and potential errors in the EHR data did not lead to spurious conclusions.
3.3.1 Preliminary Estimates Based on the Pilot and Wave I Validation
Incorporating the Pilot and Wave I, we re-fit the model with the first $52$ patients’ validated data and all $1000$ patients’ unvalidated data. Simulations in Section 2.6.1 were designed from these “preliminary SMLEs.” For an 18-year-old patient with an ALI of 0, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$). The association between ALI and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase (OR =1.21=1.21, 95% CI: $1.17,1.25$), adjusting for age. The adjusted OR =1.10=1.10 for age was relatively unchanged (OR =1.10=1.10, 95% CI: $1.01,1.21$). Preliminary SMLEs reaffirmed that worse whole-person health was associated with higher expected odds of engaging in the healthcare system, perhaps more strongly than the naive model indicated.
@@ 3.3.2 Final Estimates Based on All Waves of Validation
After completing all three waves, we used the $100$ patients’ validated data with all $1000$ patients’ unvalidated data to re-fit the SMLEs one last time. These final estimates all fell between the naive model and preliminary SMLEs. The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 0 were $0.23$ (95% CI: $0.16$, $0.33$). For each $0.1$-point increase in ALI, a $12\%$ increase in odds was expected (OR =1.12=1.12, 95% CI: $1.09,1.15$), adjusting for age. Again, older patients were expected to be more likely to engage (adjusted OR =1.11=1.11, 95% CI: $1.02,1.20$). Technical details and R code are in the Supplemental Materials.
@@ 3.4 Informing Implementation in the Learning Health System
All models agreed that worse whole-person health (higher ALI) was associated with a higher likelihood of engaging in the healthcare system, adjusting for age. Chart reviews uncovered high quality of non-missing EHR data and located auxiliary information for some missing data. Without validation, knowing whether we could trust the EHR data and whether the EHR-derived ALI was a valid computable phenotype for whole-person health would not have been possible. Thus, while the models led to the same conclusions, validating EHR data before using them in practice or research is critical.
Our findings can now inform implementation strategies to incorporate the ALI into this learning health system. For example, if the ALI automatically prepopulates in patients’ charts, clinicians would have easy access to it when making care decisions, and researchers could extract it for further studies. This healthcare system previously implemented another score, the EFI, across the patient population aged $65$ and older. Building on that experience and encouraged by the ALI’s predictive ability, we plan to operationalize it for patients aged $18$ and older.
Of course, we want to embed the validated ALI, and only a small subset of patients can undergo chart reviews. Fortunately, in obtaining SMLEs for the healthcare utilization model, we robustly estimate the exposure error mechanism, $\Pr(X|X^{*},Z)$. When scaling up across the healthcare system, the estimated exposure error mechanism can be used to efficiently predict patients’ validated ALI from unvalidated EHR data (Figure 7). See the Supplemental Materials for details on calculating these predictions.
Figure 7: Among patients with chart reviews (triangles), validated allostatic load index (ALI) could differ noticeably from the unvalidated version using the unvalidated electronic health records (EHR) data. For patients without chart reviews (squares), validated ALI could be robustly predicted using the estimated exposure error mechanism. The dashed line indicates the line of equality (i.e., where unvalidated and validated ALI are the same). The solid line is a loess smoother.
@@ 4 Discussion
Predicting healthcare utilization is a priority in the learning health system, with important implications at both patient and institutional levels. Establishing standardized, valid measures of whole-person health (like the ALI) that are associated with utilization and can scale up to the entire institution is an integral step. While EHR data contain variables to calculate these measures, missing data is a key challenge, and non-missing data are still prone to errors.
@@
Data validation, often through chart reviews, can gauge data quality, but this time-intensive process is usually only possible for a small subset of patients. Also, existing validation protocols do not overcome missingness. We propose novel methods to maximize information gained from a partial validation study through the statistical model, validation protocol, and targeted patient selection strategy. We focused on estimating the association between ALI and healthcare utilization. In preparation for EHR implementation, partial validation can also efficiently improve estimates of the ALI itself.
@@
The proposed analytical and design strategies were devised to ensure robust and efficient estimates. Specifically, the statistical model avoids making restrictive assumptions about the error mechanism, while incorporating all available information to achieve precise estimates (i.e., low variance). Moreover, the targeted design strategies are tailored to the preliminary validation data and chosen statistical model, selecting patients who are most informative for the clinical question. Our enriched validation protocol builds on the literature (e.g., [52]) by incorporating a clinically-driven “roadmap” to locate auxiliary information about missing values.
@@
Due to the curse of dimensionality, the SMLEs can handle conditioning on error-prone $X^{*}$ and maybe one other covariate $Z$ in the error mechanism. We assumed that validated ALI depended only on its unvalidated version. However, if a more complicated error structure is expected, data pre-processing will be needed before using the SMLEs [53].
@@
While our roadmap recovered some missing ALI components, validated patients could still have incomplete data. Clinical guidelines were only defined for missing values presumed to be unhealthy. Missingness could be further reduced if auditors were given additional guidance on auxiliary information for healthy values. Others have described missingness in the EHR as selection bias and corrected for it using approaches like IPW [54]. Applying such approaches to our post-validation data may reduce the impact of residual missingness. Alternatively, we could adopt other allostatic load calculations, like the index of cardiometabolic health, which rely less on rare labs and may face fewer missing components [8].
@@
EHR data are susceptible to “informative presence bias,” wherein the process through which patients engage in the healthcare system can bias associations [55, 56, 57]. Our decision to adapt the Seeman et al. (2001) [9] ALI as the proportion of non-missing components could be considered a summary measure. A recent review found summary measures to be the most common approach to handling informative presence bias [58].
@@
The ALI is among many computable phenotypes currently applied to EHR data [16]. Other phenotyping algorithms, including EHRshot [59], translate observable characteristics into measures of underlying patient states. Validation is a critical step in their development and typically involves chart reviews to obtain “gold standard” labels [60, 61]. When phenotyping with machine learning methods, this validation set generally needs to be an SRS. Meanwhile, our statistical model accommodates targeted validation sampling, creating opportunities to promote the efficiency of downstream analyses (e.g., associations between the phenotype and suspected risk factors).
@@
There are several interesting directions for future work. First, a longitudinal approach to modeling healthcare utilization, wherein patients’ ALI can vary over time, might provide more information. Second, auditors recorded their findings in Microsoft Excel, but an electronic REDCap [62] interface would improve usability and accuracy [34]. Third, targeted validation strategies could be used to improve predictive ability rather than efficiency [63, 64].
@@ 5 Conclusion
EHR data provide an exciting opportunity to implement key measures in the learning health system. However, errors and missingness in these data pose challenges in verifying and operationalizing such measures. Validation of even a subset of patients’ data is time- and cost-intensive but provides crucial insights into data quality. Statistical and biomedical informatics methods can maximize information gained from such partial validation.